OpenStudy (anonymous):

Write the equation of the line in point-slope form: given (-3, -2) and the slope of 1

OpenStudy (anonymous):

Point-slope form is: \[y-y _{1} = m(x - x _{1})\] Substitute the values of (-3, -2) in the formula on the variables \[x _{1}, y_{1}\] \[y - (-2) = (1)(x - (-3))\] \[y +2 = 1(x + 3)\] or simply as \[y +2 = x + 3\] Voila!

OpenStudy (precal):